5. Which category of inventory holding costs would be much higher than average for rapidly changing industries such as pesonal computers and cell phones?

a Obsolescence

b. Warehouse Leasing

c. Labor handling

 

6. High stock-out costs relative to holding costs should lead to recommendations for larger orders

a. Yes b. No

 

7. I dhe oruantity (ize) doubles, what happensto the mumber of onders placed per year? (hint: # of orders)T-Demand EOQ)

a. Declines by more than 50%

b, Declines by 50%

c. Increases by 50%

d. Does not change .

 

8. Instead of following the EOQ prescription of 150 pieces/order, a company orders 200 pieces at a time. This action would typically result in:

a. Higher total ordering cost, but lower total holding cost

b. Lower total ordering cost, but higher total holding cost

c. Lower total ordering cost and lower total holding cost

d. Higher total ordering cost and higher total holding cost
Answer #1

5. obsolescence has a much higher cost as to either store or to dispose the inventory there is a high cost.

6. Yes. With larger orders, there is a higher customer service level. If stockout costs> holding cost, it means we are placing smaller orders compared to demand.

7. Let's assume total demand as 1000 and order size as 100. Total orders = 1000/100 = 10. When order size doubles, 200 . Number of orders = 1000/ 2000 = 5. Hence, declines by 50%.

8. Option b. Holding cost will increase as average inventory will increase with larger orders. However, ordering cost will reduce as the number of orders placed will be lesser.